OpenStudy (anonymous):

how do i evaluate (2^3)*(3^4)?

OpenStudy (anonymous):

= 2x2x2x3x3x3x3 = 8 x 81

OpenStudy (anonymous):

maybe i should ask how i multiply these exponates as they have different bases.

OpenStudy (saifoo.khan):

\[\Large 2^3 = 2 \times 2 \times 2\] \[\Large 3^4 = 3 \times 3 \times 3 \times 3\]

OpenStudy (saifoo.khan):

U can expand it.